how do u find the third, fifth, and tenth termsof the sequence described by each rule.
A(n)=4+(n+1)(3)

Answer by rfer(16322)

A(n)=4+(3+1)(3)
A(n)=4+4*3
A(n)=16
------------------
A(n)=4+(5+1)(3)
A(n)=4+6*3
A(n)=22
you try the other one.
Bob